# Bayesian sample size calculations for external validation studies of risk prediction models

### v1

@@ 2.4.2 Specifying $P(\theta)$
The first condition guarantees that the c-statistic relating $\pi$ to $Y$ (which is often reported) is equal to the c-statistic for $P(p)$ (as c-statistic is invariant under monotonical transformation of predictor values). The second condition is a requirement for the identifiability of $P(p)$ given its mean (prevalence) and c-statistic23.
@@
As an example of this identifiability, consider an outcome prevalence of 0.25, c-statistic of 0.75. calibration slope of 1.1, and O/E ratio of 0.9. Assuming predicted risks have a Logit-normal distribution, the calibrated risks will also have Logit-normal distribution. The parameters of the latter are uniquely identifiable from $\phi,c$, which resolves to $P(p)\sim\mbox{Logitnorm}(-1.3302,1.0395)$ (the mcmapper R package implements our proposed numerical algorithms for this mapping24). As well, there is a 1:1 mapping between the O/E ratio and calibration intercept. Given the calibration slope of of 1.1 and the specified distribution for $p$, an O/E ratio of 0.9 uniquely maps to calibration intercept of -0.089 (see footnote of Table 1 for additional informaiton). Thus, to generate a random validation sample, one can sample $N$ calibrated risk from $P(p)$, generate corresponding response values as $Y_{i}\sim\mbox{Bernoulli}(p_{i})$ and compute predicted risks as $\pi_{i}=h^{-1}(p_{i})$. $(\pi_{i},Y_{i})$s created this way will be a realization of $D_{N}$.
@@
Given this identifiability, characterizing our prior information involves specifying the joint distribution $P(\phi,c,h)$. Ideally, the previous analysis would report both the value and an estimated covariance matrix variance of $\hat{\theta}$, i.e., there would be joint inference about the three elements of theta. These would then become the prior mean and covariance matrix of the joint prior for $\theta$. In typical practice, however, joint inference on such parameters is not typically reported, but probability distributions for individual components can readily be constructed from existing information. For example, for outcome prevalence, if from a previous study with sample size of $n$, $m$ individuals experience the outcome, our knowledge can be specified as $\mbox{Beta}(m,n-m)$. For other components, point estimates and reported CI widths can be used, along an assumed distribution type, to construct distributions. Examples include specifying a Log-normal distribution for O/E ratio, Normal distribution for calibration slope, and Beta for c-statistic based on reported point estimates and bounds of 95%CI. The accompanying software provides a flexible way of specifying such distributions, accepting both direct specification of distribution parameters, moments, or mean and upper bound of 95%CI.
@@
Once marginal distributions are constructed, a simple strategy would be to complete the prior specification by imposing a priori independence between the three components. An optional strategy, which would be more faithful to the true data generating mechanism, would involve a parametric bootstrap procedure to recover parameters interdependence. In this approach, one simulates multiple samples given $theta=\hat{thata}$, and for each sample record the ensuing estimates of $\hat{\theta}$. The empirical correlation matrix of these simulated estimates can then be taken as the prior correlation matrix, instead of simply presuming an uncorrelated prior. Appendix 1 provides an algorithmic description of this approach.
@@
This algorithm in itself quantifies the uncertainty for a population that is exchangeable with the population(s) from which evidence on model performance is collected. If the evidence is collected from a single population, this specification assumes model performance is the same between the source and target populations. On the other hand, if current evidence is synthesized from multiple populations using meta-analytic techniques (as in our case study below), this specification assumes the target population is a random draw from the distribution of the meta-population. The predictive distribution of model performance in a new population can thus be used to characterize uncertainties. If the exchangability assumption does not hold, different steps of this algorithm can be modified to model population differences. If the outcome prevalence are expected to be different, one can shift the distribution of calibrated risks once they are determined in Step 4a to match the prevalence in the target population (which itself should be a random variable indicating our uncertainty about prevalence). Independently, if evidence is extracted from a model development study, and there are concerns about the model being overfitted, one can add a negative penalty term to the calibration slope in Step 4b (itself a random variable) representing our knowledge about the degree of overfitting. A structured examination of the degree of relatedness between the target population and source population(s) might help the investigator decide on the need for, and extent of, modifications25.
2.4.3 Sampling from posterior distributions
@@
All sample size rules require expectation with respect to the distribution of parameters ($\theta$) and future sample ($D_{N}$). This naturally invites a Monte Carlo sampling algorithm based on repeated sampling from $P(\theta)$, simulating validation samples from $P(D_{N}|\theta)$, and computing the precision targets and VoI metrics in this sample. Draws from the posterior distributions are then processed according to the sample size rules of interest (e.g., average CI widths for the expected CI width criterion, the corresponding quantile for assurance-based metrics, or VoI for NB).
@@
In addition to this default approach based on simulating $D_{N}$, we propose an approximate two-level closed-form approach for CI width targets that does not involve simulating $D_{N}$. This approaches is based on using SE equations twice. Let $\theta_{k}$ be the metric of interest (e.g., c-statistic) among $\theta$s. First, in the jth iteration of the Monte Carlo simulation, we specify $P(\hat{\theta}_{k}|\theta)$, the sample distribution of the metric of interest in the validation sample given our draw from $\theta$. This distribution is specified via method of moments with the first two moments being the corresponding value for $\theta_{k}$ in $\theta$, and $SE^{2}(\theta_{k})$ from the relevant SE equation, which is interpreted in a Bayesian flavor as the distribution of the future sample estimate. We then obtain a draw from such a distribution as a realization of $\hat{\theta}_{k}$, which is plugged into the SE equation again to quantify its CI width.
@@
Table 1: Bayesian Monte Carlo algorithm for drawing from the posterior distribution of precision targets
1. Assign a distribution to prevalence, c-statistic, and calibration function representing current knowledge (e.g., based on reported point estimates and confidence bands from previous development or validation studies). If calibration function is specified as a line on the logit scale, this can be one of the following. Distributions for calibration intercept and slope. Distributions for O/E ratio and calibration slope. Distributions for mean calibration and calibration slope. 2. Assign a distribution type for calibrated risks $p$ in the source population 3. Obtain a sample of size S for $\theta$: $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\},j=1,2,...,S$. Optional: use parametric bootstrapping to induce correlation among $\{\phi,c,h\}$ 4. For j=1 to S (number of Monte Carlo simulations). (a) Derive the parameters of $P^{(j)}(p)$, the distribution of calibrated risks in this iteration, given $\phi^{(j)}$ and $c^{(j)}$, given the distribution type assigned in step 2. *Sample-based method (b) Draw $N$ observations for calibrated risks: $p_{i}^{(j)}\sim P^{(j)}(p),i=1,2,...,N$. Draw corresponding response values $P(Y_{i}^{(j)})\sim\mbox{Bernoulli}(p_{i}^{(j)})$. Calculate the corresponding values of $\pi$: $\pi_{i}^{(j)}={h^{(j)}}^{-1}(p_{i}^{(j)})$. Construct ${D_{N}}^{(j)}=\{(\pi_{i}^{(j)},Y_{i}^{(j)})\}_{i=1}^{N}$ as the validation sample. (c) Using ${D_{N}}^{(j)}$, construct and record precision targets (CI widths): $ciw^{(j)}=g({D_{N}}^{(j)})$ for each metric of interest. *Two-step closed-form method (b) For any metric $\theta_{k}$, specify $P(\hat{\theta}_{k}|\theta^{(j)})$ using method of moments, with the first moment being true value of $\theta_{k}$ from $\theta^{(j)}$, and the second moment being $SE^{2}(\theta_{k})$ (and a choice of distribution type#). Obtain $\hat{\theta}_{k}$ as a draw from this distribution. (c) Plug $\hat{\theta}_{k}$ into the relevant SE equation to compute the precision criterion (CI width). Record this value. Next j 5. Process the draws from the precision targets according the sample size rule specified:
@@
For VoI analysis, we are not aware of any previous work proposing optimality assurance. As well, the previously proposed algorithms for validation EVSI were fully Bayesian18. Our proposed algorithm for computing optimality assurance and EVSI for the mixed Bayesian-likelihood setup is presented in Table 2. We note that given that sample estimates of prevalence, sensitivity, and specificity all depend on the four frequencies $\{N_{tp},N_{fn},N_{tn},N_{fp}\}$ (the number of true positives, false negatives, true negatives, and false positives, respectively), $D_{N}$ can be fully specified by these four frequencies, which in turn can be drawn from their respective binomial distributions without simulating individual-level data.
@@
(e) Calculate sample esaimtes of NBs: ${NB}_{D_{N}^{(j)}}(k);k\in\{0,1,2\}$, by plugging in the sample estimates of prevalence, sensitivity, and specificity from ${D_{N}}^{(j)}$ in equation 2.2. (f) Identify the winning strategy in the sample: $W^{(j)}=\mbox{argmax}_{k}{NB}_{D_{N}^{(j)}}(k)$. (g) Record the true NB of this strategy $NBsample^{(j)}={NB}_{\theta}(W^{(j)})$. (h) Record whether the winning strategy had the highest possible NB: $A^{(j)}=I({NB}_{\theta}(W^{(j)})=NBmax^{(j)})$
@@ 2.4.4 Implementation
The above algorithms are implemented in the bayespmtools package (https://github.com/resplab/bayespmtools), in particular as two functions bpm_valprec() and bpm_valsamp(), for, respectively, computing precision / VoI for a fixed sample size, or determining the sample size corresponding to a set of precision / VoI criteria and sample size rules. As input, these functions expect evidence to be parameterized via four probability distributions, one for outcome prevalence, c-statistic, calibration slope, and one of calibration intercept, O/E ratio, or mean calibration. Missing correlation among these distributions is optionally (active by default) imputed via the parametric bootstrap method explained earlier. The correlation is induced to the sample of draws from marginal draws for these parameters via the methods by Iman and Conover26. This method offers the flexibility to the investigator to use different distribution types for each component (as opposed to a single multivariate distribution). Both algorithms currently implement Logit-normal, Probit-normal, or Beta for the distribution of calibrated risks.
@@
bayespm_prec() computes expected CI widths, quantiles of CI widths, as well as optimality assurance and EVSI using the above-mentioned Monte Carlo sampling algorithms. bayespm_samp() solves for minimum $N$ that satisfies any of the requested precision criteria coupled with a sample size rule (e.g., targeting an expected CI width of 0.2 for O/E ratio, or 90% assurance that c-statistic CI width<0.1). Because for any $N$, one iteration of the Monte Carlo simulation generates one draw from the posterior distribution of CI widths or NBs, finding the minimum $N$ that satisfies a given sample size rule is a stochastic root-finding problem. In our implementation, for targets related to CI widths (expected value and assurance), we use the Robbins-Monro algorithm27, This was motivated by the simplicity of this algorithm which enables simultaneous optimization for all widths targets. For VoI and optimality assurance, we note that binomial draws can be vectorized across the entire sample of $\theta$s, thus these calculations can be performed faster than those based on CI Widths. We use the Stochastic-Simultaneous Optimistic Optimization algorithm by Valko et al28 (implemented in the OOR R package29).
3 Case study: Predicting deterioration in hospitalized COVID-19 patients
@@
The ISARIC 4C model is a risk prediction model for predicting deterioration (need for ventilatory support or critical care, or death) in patients hospitalized due to COVID-19 infection. Details of the development and validation on this model are provided in its original publication30. In summary, the investigators used electronic hospital records from all 9 health regions of the UK to develop and validate . The London region was left out for external validation, and the remaining 8 regions were used in internal-external validation31. One at a time, one region was left out, the model was developed in the remaining 7 regions, and its performance was evaluated in the left-out region. Random-effects meta-analysis was used to pool out-of-sample estimates of discrimination and calibration metrics. Finally, the investigators used the data from all 8 regions to fit one final model, and evaluated its performance in the London region.
@@
To use this setup as an informative example, we assume the London region did not participate in this study, and is now interested in conducting a validation study of this model in their region. That is, we ignore the results reported on the performance of the model in the London region, and consider the internal-external validation results as the information at hand. We assume that the performance of the model in the London region can be assumed to be a random draw from the distribution of performance observed across other regions (i.e., it is exchangeable with other regions).
@@
The total development sample size was 70,349 (after removing those with unknown outcome status), with total number of events (deterioration) being 30,316, giving rise to an overall outcome prevalence of 0.428. Pooled estimates of the c-statistic, mean calibration, and calibration slope were 0.76 (95%CI 0.75–0.77), -0.01 (95%CI -0.12–0.09), and 0.99 (95%CI 0.97–1.02), respectively.
3.0.1 Application of Riley approach
@@
As a baseline comparison, we performed sample size calculations for validating this model using the multi-criteria method proposed by Riley et al3. Riley et al used the ISARIC 4C as an example, and for consistency, we use the same targets: confidence interval widths of 0.1 for the c-statistic, 0.22 for O/E ratio, 0.3 for calibration slope. The assumed true values of these metrics was taken as the point estimates from the above-mentioned meta-analyses. The results indicate a sample size of 1,056 is required, which is dictated by the calibration slope. For other components, sample sizes are as follows: O/E ratio: 425, c-statistic: 359.
3.0.2 Application of Bayesian approach
@@
Specifying $P(\theta)$: Given our assumption about the exchangeability of populations across UK geographic regions, the predictive distribution of the model performance in a new region represents our uncertainty about the model performance in the London region. However, it is crucial to note that pooled estimates and confidence bands reported in the original study are for the average effect and do not represent predictive distributions. To obtain predictive distributions, we re-performed the meta-analyses pooling c-statistic, mean calibration (note that while we specify evidence in terms of mean calibration, as is reported by Gupta et al, we target O/E ratio for sample size determination), and calibration slope for the 8 regions in the internal-external validation (we extracted study-specific estimates from digitized figures which enables us to use three significant digits). Gupta et al did not perform a meta-analysis on prevalence. We therefore performed this additional meta-analysis using region-specific prevalence data from their report. Given the large sample sizes, we performed all meta-analyses based on normality assumption on the original scale of each parameter, aside from c-statistic, for which a logit-transformation was performed as recommended by Snell et al32. We used method of moments to recover distribution from the mean and SD of the predictive distribution for a new population. We specified Beta distributions for prevalence and c-statistic, and Normal distributions for mean calibration and calibration slope. This resulted in the distributions for the four parameters presented in Table LABEL:tab:ISARIC_evidence.
@@
Parameter Distribution Mean..SD. CI.bounds
@@
There are some approximations in the above approach that are made in favor of practicality. Possible correlations between parameters within each study were not available.We did not model between-study correlations (e.g., via a joint meta-analysis of parameters), after an exploratory analysis (examining Spearman rank correlation coefficients) did not support the presence of strong between-study correlation.
Target precision / VoI criteria and sample size rules: For comparability, we target the same CI width as in the Riley approach, targeting 95%CI width of, 0.22 for O/E ratio, 0.10 for c-statistic, and 0.30 for calibration slopes. For these metrics, we apply both the expected value and assurance rules. That is, we determine sample sizes that correspond to the expected CI width being equal to our targets (this can be interpreted as a direct Bayesian counterpart of conventional Riley et al’s formula). We also demand a 90% assurance to meet or exceed the CI width criteria, representing our preference for obtaining a narrower over wider CI width than the targeted with. As for clinical utility, we demand a 90% assurance on NB at the 0.2 threshold; that is, we desire to be 90% confident that the strategy that will emerge as the best in the validation sample is actually the best strategy in the target population. Finally, we investigate the expected gain in NB for each component of the sample size on the EVSI curve.
Analysis setup: All analyses are based on 10,000 draws from $P(\theta)$, assuming Logit-normal distribution for calibrated risks. This assumption is tested in a sensitivity analysis. We used the sample-based approach in this analysis. Results of the two-level closed-form approach are provided in Appendix 2.
Results: Table 4 provides the sample size for each component. The expected CIW components are close to their frequentist counterparts. The assurance-based targets result in slightly higher sample sizes, as expected. The final sample size of 1173 is dictated by the assurance component of the calibration slope.
@@
Bayesian (expected CI width) 360 426 1025
Bayesian (90% assurance) 397 516 1173 306
Figure 1 shows how the precision targets change with sample size. These are computed from a separate Monte Carlo simulation of size 10,000 (independently of the calculations that determined the sample sizes) and as such act as diagnostic plots of CIW and NB assurance.
Figure 1: Expected CI widths (a) 90th quantile of CI width (b), optimality assurance (c) curves. Shapes on the line pertain to individual sample size components.
@@
(a) Expected value of CI width (b) 90th percentile of CI widths (c) Optimality assurance
Figure 2 shows the exemplary kernel histograms of the distribution of CI widths (the first three panels) for the smallest ($N=306)$ and largest ($N=1,173$) components of the sample size. The last panel demonstrates the distribution of the incremental NB of the model compared with the default strategies (i.e., $NB_{1}-\max(NB_{0}-NB_{2})$). With higher sample sizes, the CI widths get both shorter and also get more clustered. The distributions are relatively symmetrical. This can explain why the eciw values are close to their conventional, frequentist counterparts. For NB, as the sample estimate of NB is an unbiased estimator, higher sample sizes will result in a narrower distributions. The assurance for NB, and for EVSI, what affects the results is the share of the distribution that falls on the left side of zero.
@@
Turning to VoI analysis, Figure 3 provides the EVSI calculation (with EVSI/EVPI ratio as the secondary Y-axis). Components of sample size calculation are overlaid on the graph.

### v2

@@ Methods
The first condition guarantees that the c-statistic relating $\pi$ to $Y$ (which is often reported) is equal to the c-statistic for $P(p)$ (as c-statistic is invariant under monotonical transformation of predictor values). The second condition is a requirement for the identifiability of $P(p)$ given its mean (prevalence) and c-statistic[24].
@@
As an example of this identifiability, consider an outcome prevalence of 0.25, c-statistic of 0.75, calibration slope of 1.1, and O/E ratio of 0.9. Assuming predicted risks have a Logit-normal distribution, the calibrated risks will also have Logit-normal distribution. The parameters of the latter are uniquely identifiable from $\{\phi,c\}$, which resolves to $P(p)\sim\mbox{Logitnorm}(-1.3302,1.0395)$ (the mcmapper R package implements our proposed numerical algorithms for this mapping[25]). As well, there is a 1:1 mapping between the O/E ratio and calibration intercept. Given the calibration slope of of 1.1 and the specified distribution for $p$, an O/E ratio of 0.9 uniquely maps to calibration intercept of -0.089 (see footnote of Table 1 for additional informaiton). Thus, to generate a random validation sample, one can sample $N$ calibrated risk from $P(p)$, generate corresponding response values as $Y_{i}\sim\mbox{Bernoulli}(p_{i})$ and compute predicted risks as $\pi_{i}=h^{-1}(p_{i})$. $(\pi_{i},Y_{i})$s created this way will be a realization of $D_{N}$.
@@
Given this identifiability, characterizing our prior information involves specifying the joint distribution $P(\phi,c,h)$. Ideally, the previous analysis would report both the value and an estimated covariance matrix variance of $\hat{\theta}$; i.e., there would be joint inference about the three elements of $\theta$. These would then become the prior mean and covariance matrix of the joint prior for $\theta$. In typical practice, however, joint inference on such parameters is not reported, but probability distributions for individual components can readily be constructed from existing information. For example, for outcome prevalence, if from a previous study with sample size of $n$, $m$ individuals experience the outcome, our knowledge can be specified as $\mbox{Beta}(m,n-m)$. For other components, point estimates and reported CI bounds can be used, along an assumed distribution type, to construct distributions. Examples include specifying a Log-normal distribution for O/E ratio, Normal distribution for calibration slope, and Beta distribution for c-statistic based on reported point estimates and bounds of 95%CI. The accompanying software provides a flexible way of specifying such distributions, accepting specification of distribution parameters, moments, or mean and upper bound of 95%CI.
@@
Once marginal distributions are constructed, a simple strategy would be to complete the prior specification by imposing a priori independence between the three components. An optional strategy, which would be more faithful to the true data generating mechanism, would involve a parametric bootstrap procedure to recover parameters interdependence. In this approach, one simulates multiple samples given $\theta=\hat{\theta}$, and for each sample records the ensuing estimates of $\hat{\theta}$. The empirical correlation matrix of these simulated estimates can then be taken as the prior correlation matrix, instead of simply presuming an uncorrelated prior.
@@
This algorithm in itself quantifies the uncertainty for a population that is exchangeable with the population(s) from which evidence on model performance is collected. If the evidence is collected from a single population, this specification assumes model performance is the same between the source and target populations. On the other hand, if current evidence is synthesized from multiple populations using meta-analytic techniques (as in our case study below), this specification assumes the target population is a random draw from the distribution of the meta-population. The predictive distribution of model performance in a new population can thus be used to characterize uncertainties. If the exchangability assumption does not hold, different steps of this algorithm can be modified to model population differences. If the outcome prevalence are expected to be different, one can shift the distribution of calibrated risks once they are determined in Step 4a to match the prevalence in the target population (which itself should be a random variable indicating our uncertainty about prevalence). Independently, if evidence is extracted from a model development study, and there are concerns about the model being overfitted, one can add a negative penalty term to the calibration slope in Step 4b (itself a random variable) representing our knowledge about the degree of overfitting. A structured examination of the degree of relatedness between the target population and source population(s) might help the investigator decide on the need for, and extent of, modifications[26].
Sampling from pre-posterior distributions
@@
All sample size rules require expectation with respect to the distribution of parameters ($\theta$) and future sample ($D_{N}$). This naturally invites a Monte Carlo sampling algorithm based on repeated sampling from $P(\theta)$, simulating validation samples from $P(D_{N}|\theta)$, and computing the precision targets and VoI metrics in this sample. Draws from the pre-posterior distributions are then processed according to the sample size rules of interest (e.g., average CI widths for the expected CI width criterion, the corresponding quantile for assurance-based metrics, or VoI for NB).
@@
In addition to this default approach based on simulating $D_{N}$, we propose an approximate two-step approach for CI width targets that does not involve simulating $D_{N}$. This approaches is based on using SE equations twice. Let $\theta_{k}$ be the metric of interest (e.g., c-statistic) among $\theta$s. First, in the jth iteration of the Monte Carlo simulation, we specify $P(\hat{\theta}_{k}|\theta)$, the sample distribution of the metric of interest in the validation sample given our draw from $\theta$. This distribution is specified via method of moments with the first two moments being the corresponding value for $\theta_{k}$ in $\theta$, and $SE^{2}(\theta_{k})$ from the relevant SE equation, which is interpreted in a Bayesian flavor as the distribution of the future sample estimate. We then obtain a draw from such a distribution as a realization of $\hat{\theta}_{k}$, which is plugged into the SE equation again to quantify its CI width.
@@
Table 1: Bayesian Monte Carlo algorithm for drawing from the pre-posterior distribution of precision targets
1. Assign a distribution to prevalence, c-statistic, and calibration function representing current knowledge (e.g., based on reported point estimates and confidence bands from previous development or validation studies). If calibration function is specified as a line on the logit scale, this can be one of the following. Distributions for calibration intercept and slope. Distributions for O/E ratio and calibration slope. Distributions for mean calibration and calibration slope. 2. Assign a distribution type for calibrated risks $p$ in the source population 3. Obtain a sample of size S for $\theta$: $\theta^{(j)}=\{\phi^{(j)},c^{(j)},h^{(j)}\},j=1,2,...,S$. Optional: use parametric bootstrapping to induce correlation among $\{\phi,c,h\}$ 4. For j=1 to S (number of Monte Carlo simulations). (a) Derive the parameters of $P^{(j)}(p)$, the distribution of calibrated risks in this iteration, given $\phi^{(j)}$ and $c^{(j)}$, given the distribution type assigned in step 2. *Sample-based method (b) Draw $N$ observations for calibrated risks: $p_{i}^{(j)}\sim P^{(j)}(p),i=1,2,...,N$. Draw corresponding response values $P(Y_{i}^{(j)})\sim\mbox{Bernoulli}(p_{i}^{(j)})$. Calculate the corresponding values of $\pi$: $\pi_{i}^{(j)}={h^{(j)}}^{-1}(p_{i}^{(j)})$. Construct ${D_{N}}^{(j)}=\{(\pi_{i}^{(j)},Y_{i}^{(j)})\}_{i=1}^{N}$ as the validation sample. (c) Using ${D_{N}}^{(j)}$, construct and record precision targets (CI widths): $ciw^{(j)}=g({D_{N}}^{(j)})$ for each metric of interest. *Two-step approach (b) For any metric $\theta_{k}$, specify $P(\hat{\theta}_{k}|\theta^{(j)})$ using method of moments, with the first moment being true value of $\theta_{k}$ from $\theta^{(j)}$, and the second moment being $SE^{2}(\theta_{k})$ (and a choice of distribution type#). Obtain $\hat{\theta}_{k}$ as a draw from this distribution. (c) Plug $\hat{\theta}_{k}$ into the relevant SE equation to compute the precision criterion (CI width). Record this value. Next j 5. Process the draws from the precision targets according the sample size rule specified:
@@
For VoI analysis, we are not aware of any previous work proposing Optimality Assurance. As well, the previously proposed algorithms for validation EVSI were fully Bayesian[19]. Our proposed algorithm for computing Optimality Assurance and EVSI for the mixed Bayesian-likelihood setup is presented in Table 2. We note that given that sample estimates of prevalence, sensitivity, and specificity all depend on the four frequencies $\{N_{tp},N_{fn},N_{tn},N_{fp}\}$ (the number of true positives, false negatives, true negatives, and false positives, respectively), $D_{N}$ can be fully specified by these four frequencies, which in turn can be drawn from their respective binomial distributions without simulating individual-level data.
@@
(e) Calculate sample esaimtes of NBs: ${NB}_{D_{N}^{(j)}}(k);k\in\{0,1,2\}$, by plugging in the sample estimates of prevalence, sensitivity, and specificity from ${D_{N}}^{(j)}$ in equation Common metrics of model performance. (f) Identify the winning strategy in the sample: $W^{(j)}=\mbox{argmax}_{k}{NB}_{D_{N}^{(j)}}(k)$. (g) Record the true NB of this strategy $NBsample^{(j)}={NB}_{\theta}(W^{(j)})$. (h) Record whether the winning strategy had the highest possible NB: $A^{(j)}=I({NB}_{\theta}(W^{(j)})=NBmax^{(j)})$
@@
The above algorithms are implemented in the accompanying bayespmtools package[27], in particular as two functions bayespm_prec() and bayespm_samp(), for, respectively, computing precision / VoI for a fixed sample size, or determining the sample size corresponding to a set of precision / VoI criteria and sample size rules. These functions expect evidence to be parameterized via four probability distributions, one for outcome prevalence, c-statistic, calibration slope, and one of calibration intercept, O/E ratio, or mean calibration. Missing correlation among these distributions is optionally (active by default) imputed via the parametric bootstrap method explained earlier. The correlation is induced to marginal draws for these parameters via the methods by Iman and Conover[28]. This method offers the flexibility to use different distribution types for each component (as opposed to a single multivariate distribution). Both algorithms currently implement Logit-normal, Probit-normal, or Beta for the distribution of calibrated risks.
@@
For a specified sample size, bayespm_prec() computes expected CI widths, quantiles of CI widths, as well as Optimality Assurance and EVSI using the above-mentioned Monte Carlo sampling algorithms. bayespm_samp() solves for minimum $N$ that satisfies any of the requested precision criteria coupled with a sample size rule (e.g., targeting an expected CI width of 0.2 for O/E ratio, or 90% assurance that c-statistic CI width<0.1). Because for any $N$, one iteration of the Monte Carlo simulation generates one draw from the pre-posterior distribution of CI widths or NBs, finding the minimum $N$ that satisfies a given sample size rule is a stochastic root-finding problem. In our implementation, for targets related to CI widths (expected value and assurance), we use the Robbins-Monro algorithm[29]. This was motivated by the simplicity of this algorithm which enables simultaneous optimization for all widths targets. For Optimality Assurance and VoI metrics, the binomial draws can be vectorized across the entire sample of $\theta$s; thus these calculations can be performed faster than those based on CI widths. We use the Stochastic-Simultaneous Optimistic Optimization algorithm by Valko et al[30] (implemented in the OOR R package[31]).
Case study: Predicting deterioration in hospitalized COVID-19 patients
@@
The ISARIC 4C model is a risk prediction model for predicting deterioration (need for ventilatory support or critical care, or death) in patients hospitalized due to COVID-19 infection[32]. The investigators used electronic hospital records from all 9 health regions of the UK to develop and validate this model. The London region was left out for external validation, and the remaining 8 regions were used in internal-external validation, where one at a time, one region was left out; the model was fitted using data from all other regions, and its out-of-sample performance was assessed in the left-out region[33]. Random-effects meta-analysis was used to pool out-of-sample estimates of discrimination and calibration metrics. Finally, the investigators used the data from all 8 regions to fit one final model, and evaluated its performance in the London region.
@@
To use this setup as an informative example, we assume the London region did not participate in this study, but is now interested in conducting a validation study of this model in their region. That is, we ignore the results reported on the performance of the model in the London region, and consider the internal-external validation results as the information at hand. We assume that the performance of the model in the London region will be a random draw from the distribution of performance observed across other regions (i.e., it is exchangeable with other regions).
@@
The total development sample size was 70,349 (after removing those with unknown outcome status), with total number of events (deterioration) being 30,316, giving rise to an overall outcome prevalence of 0.428. Pooled estimates of the c-statistic, mean calibration (difference between average observed and predicted risks), and calibration slope were 0.76 (95%CI 0.75–0.77), -0.01 (95%CI -0.12–0.09), and 0.99 (95%CI 0.97–1.02), respectively. The research question now is: what (minimum) sample size is recommended for a new external validation study in the London region?
Application of Riley approach
@@
As a baseline comparison, we performed sample size calculations for validating this model using the multi-criteria method proposed by Riley et al[3]. Riley et al used the ISARIC 4C as an example, and for consistency, we use the same targets: confidence interval widths of 0.1 for the c-statistic, 0.22 for O/E ratio, and 0.3 for calibration slope. The assumed true values of these metrics were taken as the summary (pooled) estimates from the above-mentioned meta-analyses. The results indicate a sample size of 1,056 is required, which is dictated by the calibration slope. For other components, sample sizes are as follows: O/E ratio: 425, c-statistic: 359.
Application of Bayesian approach
@@
Specifying $P(\theta)$: Given our assumption about the exchangeability of populations across UK geographic regions, the predictive distribution of the model performance in a new region represents our uncertainty about the model performance in the London region. However, it is crucial to note that pooled estimates and confidence bands reported in the original study are for the average effect and do not represent predictive distributions. To obtain predictive distributions, we re-performed the meta-analyses pooling c-statistic, mean calibration (note that while we specify evidence in terms of mean calibration, as is reported by Gupta et al, we target O/E ratio for sample size determination), and calibration slope for the 8 regions in the internal-external validation. Gupta et al did not perform a meta-analysis on prevalence. We therefore performed this additional meta-analysis using region-specific prevalence data from their report. Given the large sample sizes, we performed all meta-analyses based on normality assumption on the original scale of each parameter, aside from c-statistic, for which a logit-transformation was performed as recommended by Snell et al[34]. We used method of moments to derive distribution parameters from the mean and SD of the predictive distribution for a new population. We specified Beta distribution for prevalence, Logit-normal for c-statistic, and Normal for mean calibration and calibration slope. This resulted in the distributions for the four parameters presented in Table 3.
@@
Parameter Distribution Mean(SD) 95% Credible Interval
@@
We extracted study-specific estimates from digitized figures from the original report which enables us to use three significant digits.
Matched to the 95% prediction interval from the random-effects meta-analysis - see text.
We did not model between-study correlations (e.g., via a joint meta-analysis of parameters), after an exploratory analysis (examining Spearman rank correlation coefficients) did not support the presence of strong between-study correlation.
Target precision / VoI criteria and sample size rules: For comparability, we target the same CI width as in the Riley approach, targeting 95%CI width of 0.22 for O/E ratio, 0.10 for c-statistic, and 0.30 for calibration slopes. For these metrics, we apply both the expected value and assurance rules. That is, we determine sample sizes that correspond to the expected CI width being equal to our targets (this can be interpreted as a direct Bayesian counterpart of conventional Riley et al’s targets). We also demand a 90% assurance to meet or exceed the CI width criteria, representing our preference for obtaining a narrower over wider CI width. As for clinical utility, we demand a 90% assurance on NB at the 0.2 threshold; that is, we desire to be 90% confident that the strategy that will emerge as the best in the validation sample is actually the best strategy in the target population. Finally, we investigate the expected gain in NB for each component of the sample size on the EVSI curve. This is an example of a hybrid use of this framework: CI widths and Optimality Assurance are used as sample size rules, and resulting sample sizes are assessed on in terms of their EVSI values.
Analysis setup: All analyses were performed in R[35]. We used the meta package for random-effects meta-analysis of internal-external validation results of ISARIC[36]. Sample size computations were done using the bayespmtools package, with 10,000 draws from $P(\theta)$, and assuming Logit-normal distribution for calibrated risks. We used the sample-based approach for this analysis.
Results: Table 4 provides the sample size for each component. The expected CI width components are close to their frequentist counterparts. The assurance-based targets result in slightly higher sample sizes, as expected. The final sample size of 1,181 is dictated by the assurance component of the calibration slope.
@@
Bayesian (expected CI width) 351 430 1064 NA
Bayesian (90% assurance) 399 522 1181 306
Not computed as this framework emphasizes computing assurance and VoI metrics for NB.
Figure 1 shows how the precision targets change with sample size. These are computed from a separate call to bayespm_prec() (independently of the calculations that determined the sample sizes) and as such act as diagnostic plots.
Figure 1: Expected CI widths (a), 90th quantile of CI width (b), and optimality assurance (c) curves. Shapes on the line pertain to individual sample size components.
@@
(a) Expected value of CI widths (b) 90th percentile of CI widths (c) Optimality assurance
Figure 2 shows the exemplary kernel histograms of the distribution of CI widths (the first three panels) for the smallest ($N=306)$ and largest ($N=1,181$) components of the sample size. The last panel demonstrates the distribution of the incremental NB of the model compared with the default strategies (i.e., $NB_{1}-\max(NB_{0}-NB_{2})$). With higher sample sizes, the CI widths get both shorter and also more clustered. The distributions are relatively symmetrical. This can explain why the ECIW values are close to their conventional, frequentist counterparts. For NB, as the sample estimate of NB is an unbiased estimator, higher sample sizes will result in a narrower distributions but their location remains the same. For Optimality Assurance and EVSI, what affects the results is the share of the distribution that falls on the left side of zero.
@@
Turning to VoI analysis, Figure 3 provides the EVSI calculation (with EVSI/EVPI ratio [rEVSI] as the secondary Y-axis). Components of sample size calculation are overlaid on the graph.